The mflogo Package for L^{A}TEX 2ε

Copyright (C) 1994 by Ulrik Vieth

November 28, 1995

1 Introduction

This contributed package for L^{A}TEX 2ε provides the necessary font declarations needed to access the logo font family in terms of NFSS.^{1} It also provides a little example package file that shows how to define the and logos and some appropriate font changing commands in these terms.

Using this package, there should no longer be a need to define special macros for the slanted version of these logos, and it should be possible to avoid such errors as on page 2 of The L^{A}TEX Companion where the logo appears in upright shape within an italics context of a book title.

\logofamily Once you have installed the font definition file Ulogo.fd provided here, you

\textlogo can use low-level NFSS commands to access the logo fonts in your documents, even if you do not plan to use the example package mflogo.sty. Apart from defining the and logos in terms of NFSS, that package file also provides a declarative font changing command \logofamily and a font changing command \textlogo with one argument. Perhaps these font changing commands might prove to be useful for typesetting little “teasers” such as in TTN 3;2, page 28.

1Throughout this document, the term NFSS should be understood to refer to release 2 of NFSS, i.e. the release that’s part of L^{A}TEX 2ε and that’s intended to become part of L^{A}TEX3 in the future.